# Network-aided Efficient LLM Services With Denoising-inspired Prompt Compression

### v3

@@ 3 System Model
In this section, we present the system model of wireless network-aided LLM inference services, the SLM-based prompt compression method, and the wireless transmission model with constraints on power consumption and delay.
Figure 2: System model of wireless network-aided LLM services and overview of our proposed JPPO, where user-generated long prompts are first compressed through SLM-based edge computing, then transmitted with optimized power allocation via wireless networks to the LLM server, and finally inference results are returned to users.
@@ 3.1 Wireless Network-aided LLM Services
We consider a heterogeneous wireless network where a Data Center Operator (DCO) provides LLM inference services to $N$ users with diverse task requirements, e.g., prompts. As illustrated in Fig. 2, our proposed framework consists of three key components: an SLM agent deployed at user devices or edge servers for prompt compression, a JPPO scheme for reliable wireless transmission, and a target LLM for inference service. On the user side, the SLM agent leverages its semantic understanding capability to compress prompts while preserving task-critical information. The compressed prompts are then transmitted through wireless channels with jointly optimized power allocation, and finally processed by the target LLM for inference. This framework adaptively adjusts both compression ratios and transmission power based on channel conditions and prompt characteristics to achieve high LLM service quality.
@@ 3.2 Prompt Compression
Let us formally define the prompt structure and compression process. Consider an original prompt x that consists of three components:
@@
where ${\textit{{x}}}_{\text{ins}}$ represents the instruction component, ${\textit{{x}}}_{\text{dems}}$ denotes the demonstrations or examples, and ${\textit{{x}}}_{\text{que}}$ contains the specific question or task. As depicted in Fig. 1, the token length of each component is denoted by $\mathcal{L}_{\text{ins}}$, $\mathcal{L}_{\text{dems}}$, and $\mathcal{L}_{\text{que}}$ respectively. The total token length $\mathcal{L}_{x}$ of the original prompt is given by:
@@
Our SLM-based compression mechanism generates a compressed prompt $\hat{\textit{{x}}}$ with length $\mathcal{L}_{\hat{x}}$. The compression ratio $\kappa$ is defined as:
$\kappa=\frac{\mathcal{L}_{\hat{x}}}{\mathcal{L}_{x}},\quad\kappa\in[0,1]$ (3)
@@
where $\kappa=1$ indicates no compression and smaller $\kappa$ represent higher compression rates.
We use a comprehensive fidelity metric $\mathbf{f}$ capturing three essential aspects of semantic preservation during information transmission [30] to evaluate the quality of prompt compression. The fidelity metric is composed of three key components, as shown at the bottom of Fig. 2:
  • Representation accuracy ($\mathbf{f_{1}}$), which measures how accurately the compressed prompt preserves the semantic meaning of the original prompt and maintains semantic integrity. $\mathbf{f_{1}}$ is measured by the overlap-based similarity metric based on the comparison between the representations of the original prompt and the compressed prompt. The computation of $\mathbf{f_{1}}$ is to quantify the overlap between the tokens in the original and compressed prompts, calculate the number of overlapping tokens between the two prompts, and normalize it by dividing by the length of the original prompt.
  • Transmission completeness ($\mathbf{f_{2}}$), which evaluates the integrity of information preservation during the compression process. $\mathbf{f_{2}}$ is calculated as the token retained in the compressed prompt compared to the original prompt, considering Bit Error Probability (BEP) in the wireless transmission of the user prompt. The computation $\mathbf{f_{2}}$ is obtained by multiplying the base fidelity by the channel effect. The base fidelity is calculated as the ratio of tokens retained in the compressed prompt to the number of tokens in the original prompt, adjusted by the compression ratio. Then, the channel effect is computed to reflect the impact of errors introduced by the channel.
  • Understanding accuracy ($\mathbf{f_{3}}$), which quantifies how well the receiver (target LLM) can interpret the compressed prompt correctly. $\mathbf{f_{3}}$ is to evaluate the proximity of the response to the compressed prompt received and the predefined response to the original prompt, and is calculated as the number of overlapping tokens between the expected and original responses when prompt is not compressed, normalized by the length of the expected responses.
The overall fidelity metric $\mathbf{f}$ can be defined as a weighted sum of these components:
$\mathbf{f}=\alpha_{1}\mathbf{f_{1}}+\alpha_{2}\mathbf{f_{2}}+\alpha_{3}\mathbf{f_{3}},$ (4)
@@
where $\alpha_{1}$, $\alpha_{2}$, and $\alpha_{3}$ are weight factors determining the relative importance of each fidelity component. These weights can be adjusted based on application requirements and QoS priorities.
@@ 3.3 Denoising-inspired Prompt Compression
We propose a framework for controlling the compression ratio across multiple steps, inspired by the denoising process in denoising diffusion models, e.g., Denoising Diffusion Probabilistic Models (DDPM) [18].
@@
Figure 3: The motivation of our proposed denoising-inspired prompt compression method. Part A demonstrates the process of DDPM-generating images through a denoising process. Rather than directly generating a clear image, the denoising neural network focuses on predicting noise one step at a time, using the image from the previous step to generate the current one. Part B illustrates the denoising-inspired prompt compression. Instead of directly obtaining a fully compressed prompt, SLM can also perform gradual prompt compression to reduce the degree of information that needs to be compressed in each round of processing, minimizing information loss.
@@
As shown in Part A of Fig. 3, in DDPM’s denoising process, the denoising network predicts the noise between the current state and the previous state, gradually removing noise through this process until a clean image is obtained. Compared to many AI algorithms that directly generate images, DDPM’s approach reduces training difficulty and improves image generation performance. Similarly, when we use SLM for text compression, the SLM continuously predicts the next token based on the input text to ultimately obtain the compressed prompt, as shown in Part B of Fig. 3. If we set a very high compression ratio all at once, the SLM might lose crucial information. However, if we consider denoising-inspired prompt compression, viewing the compressed prompt as a clean image and the original long prompt as a noisy state, we can enable the SLM to perform gradual prompt compression. During each compression step, we can select a relatively larger compression ratio, reducing the degree of compression, and achieve the same compression effect through multiple iterations. This way, during each compression step, the SLM needs to consider relatively less information to reduce, resulting in a lower probability of losing important information, as shown in Fig. 4 in Section 5. Although the iterative nature introduces additional SLM processing time, the system maintains its efficiency due to SLM’s swift operation—experiments in Section 5 show that each compression round only adds approximately 2% to the total LLM inference time, while the optimized compression scheme can reduce subsequent LLM inference delay by around 40%.
@@
Our method allows flexible scheduling of compression ratios while maintaining a deterministic path toward the target compression rate. Given the target compression ratio $T_{\kappa}=1/\kappa$ (e.g., $T=16$ for 16×16$\times$16 × compression), we define a compression progress variable $t\in[0,1]$, where $t=0$ represents the original text and $t=1$ represents the fully compressed text. The compression ratio at any time $t$ is governed by:
$\alpha(t)=T_{\kappa}^{-\sigma(t)}$ (5)
@@
where $\sigma(t)$ is a monotonically increasing scheduling function with $\sigma(0)=0$ and $\sigma(1)=1$. For a sequence of $M$ compression steps, we discretize $t$ into ${t_{0},t_{1},...,t_{M}}$, where $t_{0}=0$ and $t_{M}=1$. The compression ratio between consecutive steps is given by:
$\beta(i)=\frac{\alpha(t_{i})}{\alpha(t_{i-1})}=T_{\kappa}^{-[\sigma(t_{i})-\sigma(t_{i-1})]}$ (6)
@@
To enable flexible management of the compression process, We employ different scheduling methods tailored to various input prompt types, enabling flexible management of the compression process. The behavior of the compression process can be controlled through different choices of $\sigma(t)$:
  • Linear schedule: The linear schedule is defined as
@@
  • Cosine schedule: The cosine schedule is defined as
@@
  • Quadratic schedule: The quadratic schedule is defined as
@@
Note that different types of long prompts may require different optimal compression schedules. For a uniform N-step compression process with linear scheduling, the discrete time steps are given by $t_{i}=\frac{i}{M}$, where $i\in{0,1,\ldots,M}$, with each step applying a constant compression ratio of $T_{\kappa}^{-1/M}$. The total compression at any time $t$ satisfies $\alpha(t)\cdot\alpha(0)=T_{\kappa}^{-\sigma(t)}$. This formulation ensures that the initial state maintains the original length ($\alpha(0)=1$), the final state achieves the target compression ($\alpha(1)=\frac{1}{T_{\kappa}}$), and the compression path remains continuously differentiable when using smooth scheduling functions. This framework provides a principled way to design multi-step compression strategies while maintaining precise control over the compression ratio at each step.
@@ 3.4 Energy Consumption
For one user, the total energy consumption $E$ in the one-shot LLM service request process consists of two components: encoding energy $E_{e}$ and transmission energy $E_{t}$. This can be expressed as:
@@
where $P_{\text{T}}$ is the transmit power. The encoding energy consumption $E_{e}$, which represents the energy used by the SLM encoder for prompt compression, is calculated as [31]
@@
The transmission energy consumption $E_{t}$ is
@@
where $s$ represents the bit length of the compressed prompt $\hat{\textit{{x}}}$ with compression ratio $\kappa$, and $R$ is the transmission rate that can be expressed as
$R=W\log_{2}\left(1+\gamma\right)=W\log_{2}\left(1+\frac{P_{\text{T}}gd^{-\alpha}}{\lambda^{2}}\right),$ (13)
where $W$ is the bandwidth of the offloading link between the user and DCO, $\gamma$ is the Signal-to-Noise Ratio (SNR), $g$ is the Rayleigh fading coefficient (exponentially distributed with unit mean), $d$ represents the distance between user and DCO, $\alpha$ is the path-loss exponent, and $\lambda^{2}$ represents the Gaussian noise term in the Additive White Gaussian Noise (AWGN) channel.
@@ 3.5 Service Delay and Error
The total time consumption $T$ comprises three components: encoding delay in SLM and LLM, and transmission delay, expressed as:
@@
For wireless network-aided LLM services, user prompts must be uploaded to the LLM through a wireless environment. Due to potential bit errors during wireless transmission, these errors can directly affect the fidelity of the transmitted data. Therefore, it is crucial to consider the BEP to reflect the likelihood of the textural prompt being incorrectly received or decoded in the wireless system. For $n_{\rm th}$ user, the average BEP $\eta$ can be given under various modulation formats by [32]
@@
where $\zeta\left({\cdot,\cdot}\right)$ is the upper incomplete Gamma function [33, eq. (8.350.2)], $\varphi$ represents the Probability Density Function (PDF) expression for wireless channel fading. ${{{\zeta\left({{\mu_{2}},{\mu_{1}}\tau}\right)}}/{{2\zeta\left({{\mu_{2}}}\right)}}}$ is the conditional bit-error probability, ${\mu_{1}}$ and ${\mu_{2}}$ are are parameters specific to the modulation scheme, representing different combinations of modulation and detection techniques.
4 Joint Power and Prompt Optimizatiom
This section introduces the JPPO for wireless network-aided LLM services. After formulating the problem, we propose a Double Deep Q-Network (DQN) method [34] to address the joint optimization problem.
@@ 4.1 Problem Formulation
For our JPPO framework, we formulate an optimization problem that balances three key aspects: prompt compression quality, wireless transmission efficiency, and LLM service performance. The objective is to maximize the overall fidelity while satisfying power and latency constraints in the wireless network-aided LLM service system. Specifically, the joint optimization problem can be formulated as
@@
where $P_{\text{th}}$ is the maximum allowable power consumption, $T_{\text{th}}$ represents the maximum tolerable end-to-end latency, $\mathbf{f}_{\text{th}}$ defines the minimum required fidelity. The constraints are designed to ensure practical system operation. Constraint (16a) is the energy constraint that represents the total energy budget limitation at the edge device side, introducing a critical trade-off: while higher transmission power $P_{T}$ can lead to lower BEP $\eta$ and thus improved wireless transmission fidelity $\mathbf{f_{2}}$ to enhance the overall fidelity $\mathbf{f}$, the energy constraint forces a higher compression ratio $\kappa$ to reduce both the SLM/LLM inference energy cost and wireless transmission energy consumption; however, an excessively high compression ratio can result in significant information loss from the original prompt, potentially degrading both the semantic preservation fidelity $\mathbf{f_{1}}$ and LLM service quality fidelity $\mathbf{f_{3}}$. Constraint (16b) ensures the power consumption remains within the device’s power budget, Constraint (16c) guarantees that the total service latency meets real-time requirements, Constraint (16d) maintains the quality of service by enforcing a minimum threshold on fidelity. This optimization framework allows us to find the optimal balance between compression ratio and transmission power while maintaining high-quality LLM service delivery.

### v4

@@ 3 System Model
In this section, we present the system model of wireless network-aided mobile LLM services, the SLM-based prompt compression method, and the wireless transmission model with constraints on power consumption and service delay.
Figure 2: System model of wireless network-aided mobile LLM services and overview of our proposed JPPO++, where user-generated long prompts are first compressed through SLM-based edge computing, then transmitted with optimized power allocation via wireless networks to the LLM server, and finally inference results are returned to users. The illustration of the three components in the overall fidelity metric is also demonstrated: $\mathbf{f_{1}}$ represents the representation accuracy, $\mathbf{f_{2}}$ denotes the transmission completeness, and $\mathbf{f_{3}}$ is the understanding accuracy.
@@ 3.1 Wireless Network-aided LLM Services
We consider a heterogeneous wireless network where a Data Center Operator (DCO) provides LLM inference services to $N$ users with diverse task requirements, e.g., prompts. As illustrated in Fig. 2, our proposed framework consists of three key components: an SLM agent deployed at user devices or edge servers for prompt compression, a target LLM for inference service, and a JPPO++ scheme for reliable wireless transmission. On the user side, the SLM agent leverages its semantic understanding capability to compress prompts while preserving task-critical information. The compressed prompts are then transmitted through wireless channels with jointly optimized power allocation and finally processed by the target LLM for inference. This framework adaptively adjusts both compression ratios and transmission power based on channel conditions and prompt characteristics to achieve high LLM service quality.
@@ 3.2 Prompt Compression
Let us formally define the prompt structure and compression process. Consider an original prompt x that consists of three components as:
@@
where ${\textit{{x}}}_{\text{ins}}$ represents the instruction component, ${\textit{{x}}}_{\text{dems}}$ denotes the demonstrations or examples, and ${\textit{{x}}}_{\text{que}}$ contains the specific question or task. As depicted in Fig. 1, the token length of each component is denoted by $\mathcal{L}_{\text{ins}}$, $\mathcal{L}_{\text{dems}}$, and $\mathcal{L}_{\text{que}}$, respectively. The total token length $\mathcal{L}_{x}$ of the original prompt is given by:
@@
Our SLM-based compression mechanism generates a compressed prompt $\hat{\textit{{x}}}$ with length $\mathcal{L}_{\hat{x}}$. The compression ratio $\alpha$ is defined as:
$\alpha=\frac{\mathcal{L}_{\hat{x}}}{\mathcal{L}_{x}},\quad\alpha\in[0,1]$ (3)
@@
where $\alpha=1$ indicates no compression and smaller $\alpha$ represent higher compression rates.
We evaluate the quality of prompt compression using a composite fidelity metric $\mathbf{f}$ that captures three complementary aspects: semantic preservation, transmission integrity, and interpretability by the downstream LLM. These three components jointly reflect the end-to-end effectiveness of delivering and using compressed prompts in wireless network-aided LLM services, as shown in Fig. 2:
  • Representation accuracy ($f_{1}$): measures how accurately the compressed prompt preserves the semantic content of the original prompt. It reflects the degree to which key information is retained during the compression process, independent of the channel effects. This component is computed based on token-level overlap between the original prompt x and the compressed prompt $\hat{\textit{{x}}}$ as:
    $f_{1}=\frac{|\hat{\textit{{x}}}\cap\textit{{x}}|}{|\textit{{x}}|},$ (4)
    where |x^x|^$\cap\textit{{x}}|$ x | denotes the number of overlapping tokens. This metric serves as a direct measure of semantic fidelity after the compression mechanism.
  • Transmission completeness ($f_{2}$): evaluates the fraction of the compressed prompt that is successfully transmitted under practical wireless channel conditions. It captures the effect of both prompt length reduction and bit-level errors due to signal degradation. Formally, $f_{2}$ is defined as:
    $f_{2}=\frac{|\hat{\textit{{x}}}|}{|\textit{{x}}|}\cdot\left(1-\text{BEP}(\text{SNR})\right),$ (5)
    where $\text{BEP}(\cdot)$ denotes the bit error probability as a function of the Signal-to-Noise Ratio (SNR) ratio. This term models the effective information retained after wireless transmission, taking into account real-world physical-layer impairments.
  • Understanding accuracy ($f_{3}$): assesses the utility of the compressed prompt from the LLM’s perspective. It quantifies how well the transmitted prompt enables the LLM to generate an output that aligns with the expected response to the original uncompressed input. This is measured via token-level overlap between the LLM’s output $\hat{\textit{{y}}}$ and the reference original response y as:
    $f_{3}=\frac{|\hat{\textit{{y}}}\cap\textit{{y}}|}{|\textit{{y}}|}.$ (6)
    This component captures the ultimate effectiveness of the entire prompt compression and transmission pipeline in supporting the downstream task.
Together, these three fidelity components form a holistic evaluation framework. $f_{1}$ focuses on semantic preservation during compression, $f_{2}$ captures the degradation during transmission, and $f_{3}$ reflects the downstream interpretability and task utility. The overall fidelity metric $\mathbf{f}$ can be defined as a weighted sum of these components as:
$\mathbf{f}=\phi_{1}\mathbf{f_{1}}+\phi_{2}\mathbf{f_{2}}+\phi_{3}\mathbf{f_{3}},$ (7)
@@
where $\phi_{1}$, $\phi_{2}$, and $\phi_{3}$ are weight factors determining the relative importance of each fidelity component. Additionally, these weights can be adjusted based on LLM application requirements and QoS priorities.
@@ 3.3 Denoising-inspired Prompt Compression
We propose a framework for controlling the compression ratio across multiple steps, inspired by the denoising process in DDPM [19].
3.3.1 Motivation
@@
Figure 3: The motivation of our proposed denoising-inspired prompt compression method. Part A illustrates the process of generating images using DDPM through a denoising process. Rather than directly generating a clear image, the denoising neural network focuses on predicting noise one step at a time, using the image from the previous step to generate the current one. Part B illustrates the denoising-inspired prompt compression. Instead of directly obtaining a fully compressed prompt, SLM can also perform gradual prompt compression to reduce the degree of information that needs to be compressed in each round of processing, minimizing information loss.
As shown in Part A of Fig. 3, in DDPM’s denoising process, the denoising network predicts the noise between the current state and the previous state, gradually removing noise through this process until a clean image is obtained. Compared to many AI algorithms that directly generate images, DDPM’s approach reduces training difficulty and improves image generation performance.
@@
Similarly, when we use SLM for text compression, the SLM continuously predicts the next token based on the input text to ultimately obtain the compressed prompt, as shown in Part B of Fig. 3. If we set a very high compression ratio all at once, the SLM might lose crucial information. However, if we consider denoising-inspired prompt compression, where the compressed prompt is viewed as a clean image and the original long prompt is seen as a noisy state, we can enable the SLM to perform gradual compression of the prompt. During each compression step, we can select a relatively larger compression ratio, reducing the degree of compression, and achieve the same compression effect through multiple iterations. This way, during each compression step, the SLM needs to consider relatively less information to reduce, resulting in a lower probability of losing important information, as shown in Fig. 4 in Section 5.
Although the iterative nature introduces additional SLM processing time, the system maintains its efficiency due to SLM’s swift operation. Experiments in Section 5 show that each compression round only adds approximately $2\%$ to the total LLM inference time, while the optimized compression scheme can reduce subsequent LLM inference delay by around $40\%$.
3.3.2 Compression Ratios Scheduling
@@
JPPO++ allows flexible scheduling of compression ratios while maintaining a deterministic path toward the target compression rate. Given the target compression ratio $\psi_{\kappa}=1/\kappa$, e.g., $\psi=16$ for 16×16$\times$16 × compression, we define a compression progress variable as $t\in[0,1]$, where $t=0$ represents the original text and $t=1$ represents the fully compressed text. The compression ratio at any time $t$ is given by:
$\alpha(t)=\psi_{\kappa}^{-\sigma(t)},$ (8)
where $\sigma(t)$ is a monotonically increasing scheduling function with $\sigma(0)=0$ and $\sigma(1)=1$.
@@
Considering a sequence of $M$ compression steps, we can discretize $t$ into $\{t_{0},t_{1},...,t_{M}\}$, where $t_{0}=0$ and $t_{M}=1$. The compression ratio between consecutive steps is:
$\beta(i)=\frac{\alpha(t_{i})}{\alpha(t_{i-1})}=\psi_{\kappa}^{-\left(\sigma(t_{i})-\sigma(t_{i-1})\right)}.$ (9)
@@
To enable flexible management of the compression process, we employ different scheduling methods tailored to various input prompt types, thereby allowing for efficient management of the compression process. The behavior of the compression process can be controlled through different choices of $\sigma(t)$:
  • The linear schedule is defined as:
@@
  • The cosine schedule is defined as:
@@
  • The quadratic schedule is defined as:
@@
Different types of long prompts may require different optimal compression schedules to ensure efficiency. For a uniform $M$-step compression process with linear scheduling, the discrete time steps are given by $t_{i}=\frac{i}{M}$, where $i\in\{0,1,\ldots,M\}$, with each step applying a constant compression ratio of $\psi_{\kappa}^{-1/M}$. The total compression at any time $t$ satisfies $\alpha(t)\cdot\alpha\left(0\right)=\psi_{\kappa}^{-\sigma(t)}$. This formulation ensures that the initial state maintains the original length (i.e., $\alpha\left(0\right)=1$), the final state achieves the target compression (i.e., $\alpha(1)=\frac{1}{\psi_{\kappa}}$), and the compression path remains continuously differentiable when using smooth scheduling functions. This framework offers a principled approach to designing multi-step compression strategies, allowing for precise control over the compression ratio at each step.
@@ 3.4 Energy Consumption
The total energy consumption $E$ in the one-shot LLM service request process for each user consists of two components: encoding energy $E_{e}$ and transmission energy $E_{t}$, as is given by:
@@
where $P_{\text{T}}$ is the transmit power. The encoding energy consumption $E_{e}$, which represents the energy used by the SLM encoder for prompt compression, is calculated as [31]:
@@
The transmission energy consumption $E_{t}$ is given by:
@@
where $s$ represents the bit length of the compressed prompt $\hat{\textit{{x}}}$ with compression ratio $\kappa$, and $R$ is the transmission rate that can be expressed as:
$R=W\log_{2}\left(1+\gamma\right)=W\log_{2}\left(1+\frac{P_{\text{T}}gd^{-\omega}}{\lambda^{2}}\right),$ (16)
where $W$ is the bandwidth of the offloading link between the user and DCO, $\gamma$ is the SNR, $g$ is the Rayleigh fading coefficient (exponentially distributed with unit mean), $d$ represents the distance between user and DCO, $\omega$ is the path-loss exponent, and $\lambda^{2}$ represents the Gaussian noise term in the Additive White Gaussian Noise (AWGN) channel.
@@ 3.5 Service Delay and Error
The total time consumption $T$ comprises three components: encoding delay in both SLM and LLM, and transmission delay, as is given by:
@@
For wireless network-aided LLM services, user prompts must be uploaded to the LLM through a wireless environment. Due to potential bit errors during wireless transmission, these errors can directly affect the fidelity of the transmitted data. Therefore, it is crucial to consider the BEP to reflect the likelihood that the textural prompt will be incorrectly received or decoded in the wireless system. For the $n_{\rm th}$ user, the average BEP $\eta$ can be given under various modulation formats by [32]:
@@
where $\zeta\left({\cdot,\cdot}\right)$ is the upper incomplete Gamma function [33, eq. (8.350.2)], $\varphi$ represents the Probability Density Function (PDF) expression for wireless channel fading. ${{{\zeta\left({{\mu_{2}},{\mu_{1}}\tau}\right)}}/{{2\zeta\left({{\mu_{2}}}\right)}}}$ is the conditional bit-error probability, ${\mu_{1}}$ and ${\mu_{2}}$ are parameters specific to the modulation scheme, representing different combinations of modulation and detection techniques.
4 Joint Power and Prompt Optimization
This section introduces the JPPO++ for wireless network-aided LLM services. After formulating the problem, we propose a Double Deep Q-Network (DQN) method [34] to address the joint optimization problem.
@@ 4.1 Problem Formulation
For our JPPO++ framework, we formulate an optimization problem that balances three key aspects: prompt compression quality, wireless transmission efficiency, and LLM service performance. The objective is to maximize the overall fidelity while satisfying power and latency constraints in the wireless network-aided LLM service system. Specifically, the joint optimization problem can be formulated as:
@@
where $P_{\text{th}}$ is the maximum allowable power consumption, $T_{\text{th}}$ represents the maximum tolerable end-to-end latency, and $\mathbf{f}_{\text{th}}$ defines the minimum required fidelity. The constraints are designed to ensure the practical operation of the system. Constraint (19a) is the energy constraint that represents the total energy budget limitation at the edge device side, introducing a critical trade-off: while higher transmission power $P_{T}$ can lead to lower BEP $\eta$ and thus improved wireless transmission fidelity $\mathbf{f_{2}}$ to enhance the overall fidelity $\mathbf{f}$, the energy constraint forces a higher compression ratio $\kappa$ to reduce both the SLM and LLM inference energy cost and wireless transmission energy consumption. However, an excessively high compression ratio can result in significant information loss from the original prompt, potentially degrading both the semantic preservation fidelity $\mathbf{f_{1}}$ and LLM service quality fidelity $\mathbf{f_{3}}$. Constraint (19b) ensures the power consumption remains within the device’s power budget, Constraint (19c) guarantees that the total service latency meets real-time requirements, and Constraint (19d) maintains the quality of service by enforcing a minimum threshold on fidelity. This optimization framework allows us to find the optimal balance between compression ratio and transmission power while maintaining high-quality LLM service delivery.